OpenStudy (anonymous):

if the slope of the curve x^2+y^2 = 25 is -(3/4) at point (x,y) find the value of x and y. a. 3 and 4 b. 3 and 4.2 c. 2 and 3 d. 3.5 and 4

OpenStudy (mani_jha):

Find dy/dx. 2x+2y(dy/dx)=0; dy/dx = -(x/y). So -(x/y)=-(3/4) 4x-3y=0; x=(3/4)y. Put this value of x in the curve equation and get x. Then find y.0

OpenStudy (anonymous):

is that y.0? =0 then the answer is 3 and 4 because y =0 then x = 3/4